If the quadratic equation ax*2+bx+c=0 is satisfied by g,f,g*2,f*2 (where gf is not equal to zero), then the value of (g+f) can be
(A)1
(B)-1
(C)2
(D)-2

Grade:12th pass

1 Answers

VIKASH YADAV
6 years ago
In this α and β are (g) and (f) α+.β=α^2+β^2=> αβ=1.αβ=α^2β^2α+1/α=α^2+1/α^2=>α+1/α=α^2+1/α^2(α+1/α)^2-(α+1/α)-2=0α+1/α=2. Or. α+1/α=-1α=1 or. α=-1/2±√3/2So β is also same soα+β=2 or .α+β=-1 Hence answer is (b) and (c) are correct